# A Genealogy of Multi-Sensor Foundation Models in Remote Sensing

### v1

@@ 6.4. Multiple Decoders
The loss for all of these decoders is back propagated into the encoder, resulting in representations of optical imagery that capture data from other modes. While this approach learns refined optical representations and therefore, a good encoder for optical imagery, contrastive learning approaches of a similar manner would also learn encoders for the other modes of data. Depending on the downstream tasks that are required, this may be a significant limitation for MMEarth’s effectiveness.
Feature Guided MAE (FG-MAE) (Wang et al., 2023a) also explores the utilization of multiple decoders for MAE, but does so by tasking the model to reconstruct specific features of interest in the remote sensing world. MAE’s focus on pixel-level details can limit the model’s capability in understanding satellite imagery. FG-MAE explores a variety of reconstruction objectives, including Histograms of Oriented Gradients (HOG) (Dalal and Triggs, 2005), Normalized Difference Indices (NDI) such as Normalized Difference Vegetation Index (NDVI) (Huang et al., 2021), CannyEdge (Canny, 1986), and Scale-Invariant Feature Transform (SIFT) (Lowe, 2004). FG-MAE maintains two separate models: one for optical imagery and one for SAR imagery. In EuroSat (Helber et al., 2019) and BigEarthNet (Sumbul et al., 2019) classification tasks, FG-MAE’s optical model demonstrates comparable or better results to standard MAE when only reconstructing a single one of these RS features. However, performing reconstruction on both HOG and NDI results in a substantial performance boost compared to standard MAE. FG-MAE’s SAR model finds that simply reconstructing HOG yields the best results, as it tends to work well with the noisier Sentinel-1 imagery.
A sort of Venn diagram where a variety of remote sensing masked auto-encoding foundation models are compared and contrasted to highlight their similarities and differences. The models in question train on data acquired from multiple sensors or contain imagery captured by a multispectral sensor.
@@ 6.5. Orientation Awareness
Masked Angle-Aware Auto-Encoding (M3AE) (Li et al., 2024b) emerges to address this particular problem. M3AE trains on the MillionAID dataset (Long et al., 2021), an RGB dataset derived from Google Earth imagery, and introduces a scaling center crop augmentation to every image, resulting in an augmented image where an object in the center will be rotated into a different orientation from the original. This augmented image is masked then fed into an encoder-decoder framework where the decoder attempts to reconstruct the original image. In order to ensure that the model pays attention to the re-oriented object, the encoder is fed positional embeddings that include a third dimension which tracks the angle of a given patch. M3AE also introduces an optimal transport reconstruction loss to automatically assign similar image patches for each rotated crop patch; this ensures that the rotation is taken into effect from a loss standpoint.
Rotated Varied-Size Attention (RVSA) (Wang et al., 2023b) also develops a novel approach to developing orientation-agnostic representations of RS imagery, but does so through ViT self-attention compared to image augmentation. Trained on the MillionAID dataset (Long et al., 2021), RSVA shifts away from the standard multi-head self-attention (MHSA) present in traditional CV ViTs, as full attention scales quadratically to image size. Given the potential high-resolution nature of RS imagery, RSVA leverages a varied-size window-based MHSA (VSA), where images are partitioned into non-overlapping windows, then have MHSA applied (Zhang et al., 2022). In order to learn orientation, RVSA randomly rotates and re-sizes these windows before performing self-attention within a ViT. With these self-attention tweaks in effect, standard MAE is employed for reconstruction masked MillionAID images. The result is a framework capable of swiftly handling high resolution RS imagery that learns orientation-agnostic representations of RS imagery.
@@ 6.6. Wavelength Awareness
Dynamic One-For-All (DOFA) (Xiong et al., 2024) attempts to address this problem by creating a framework that can take in data from any number of different sensors. DOFA follows a hyper-network approach (Chauhan et al., 2024), where a hypernetwork takes in the wavelength details of a given band and then generates weights that are used to initialize a downstream neural network. A monochrome image of a single band is masked, then fed into that neural network to generate an encoding. That encoding is then fed through an encoder-decoder structure responsible for reconstructing the original monochrome image of that band. This framework is designed to learn bandwidth specific features via the hyper-network and shared features via the shared encoder-decoder. DOFA is trained on SAR imagery from Sentinel-1, multispectral imagery from Sentinel-2, hyperspectral imagery from enMAP, and high resolution RGB imagery from NAIP. The resulting framework achieves SOTA performance on downstream tasks such as BigEarthNet (Sumbul et al., 2019), So2Sat (Zhu et al., 2019), and EuroSat (Helber et al., 2019).
@@
A sort of Venn diagram where a variety of remote sensing masked auto-encoding foundation models are compared and contrasted to highlight their similarities and differences. The models in question all train only on RGB imagery.
@@ 7.1. Learning Representations in the Era of Climate Change
Given the rapidly shifting nature of the Earth’s climate, learned representations should be robust enough to take these changes into account without substantial retraining. Otherwise, the re-usability of these representations becomes less desirable or reliable over time, diminishing one of the main benefits of the encoder-decoder approach. When considering the variety of modalities present in RS, there is the potential to create representations that can withstand the challenges of a changing climate, but a degree of care is required. Relying too heavily on geographic locations and temporal data could very well result in representations that quickly become outdated as the world’s climate shifts. For example, a learned representation of a specific location on a coastline could cease to be relevant due to rising sea levels. Similarly, climate change has resulted in marked seasonal shifts within the past decade, so learning representations of what a given location looks like at a given time of year may wane in relevance as well.
Careful consideration is needed to either weigh temporal and geographic data in the appropriate manner or create a framework that allows for lightweight retraining of modalities that are dependent on a particular time and space, given that climate change perpetually recontextualizes a given scene in time and space. By comparison, representations based on modalities that directly observe the Earth, such as satellite imagery and text, should be more tolerant to climate change-induced drift, as their representations are tied to understanding underlying observations directly within the data itself and are less reliant on climate context. Given that many remotely sensed datasets date back decades, there is merit in creating foundation models that perform time-series forecasting over long time periods in order to effectively learn representations which take long-term trends into account. This historical data also opens the door for assessing the capability of remote sensing foundation models to quantify historically understood climate and biome changes.
@@ 7.2. Under-Tapped Modes for Geospatial Representation Learning
Much of the current research focuses on drawing parallels to the work done on CV foundation models. Given the success of foundation models in the CV domain and the sheer amount of imagery that is available, be it from satellites or other aerial sources, this is a necessary first step. However, models such as MMCPP (Chen et al., 2023) use the CLIP contrastive learning objective and rely on other modalities, such as trajectory maps, points of interest, and aerial images. Their success should serve as a somewhat cautionary tale against becoming too invested in foundation models that purely focus on satellite and/or aerial imagery. Environmental stations measure features such as groundwater, weather, or snow melt data, and provide important climate and atmospheric information at a very high temporal resolution, often recording data on a daily or even hourly basis. More human-centric data, such as social media posts, social connectedness indices (Bailey et al., 2018), and trajectory data, also provide invaluable information for understanding more societally-focused downstream tasks. Both forms of information are not observable from space, but could provide useful modes due to their different perspective and high temporal resolution. Data on rapidly changing phenomena also tends to not be observable from space, as changes can occur too quickly for a satellite to capture, due to orbit and revisit time. GAIR (Liu et al., 2025)’s usage of INR representations for satellite imagery also provides a useful starting point for ensuring that granular, ground-level data can be effectively mapped to learned representations for satellite imagery.
@@ 7.3. Performance versus Training Expenses
Given the computational costs of training transformer models, the training expense of foundation models has rapidly become an obstacle across a variety of domains, with some models now including a section on carbon footprint (Cong et al., 2023). Given that a substantial portion of downstream tasks for remote sensing pertain to combating the negative effects of climate change, there is a compelling argument for remote sensing foundation models to focus on curtailing training expenses. Massive models which create a large carbon footprint ultimately may be counterproductive from a climate change standpoint. However, the perpetual flow of remotely sensed data which dates back decades also encourages exploring the development of an incremental learning pipeline to efficiently fine-tune foundation models with up-to-date information.
@@
Aside from a carbon footprint perspective, models with huge numbers of parameters have their downstream applications somewhat limited due to computational requirements. Some models, such as MOSAIKS (Rolf et al., 2021), focus on lightweight frameworks in order to promote access to remote sensing foundation models. MOSAIKS has competitive results on a number of downstream tasks, indicating that increasing parameters within models may follow a law of diminishing returns. On the other hand, the encoder-decoder structure that is predominant in foundation models allows for repeated, diverse usage of representations on downstream tasks. Arguably, leveraging significant resources to generate high quality embeddings as a one-time cost is the most effective path moving forward, as this approach results in SOTA performance on downstream tasks with relatively few resources necessary for future fine-tuning.
@@
It is possible that distillation networks may provide a happy medium between the desire for high quality embeddings and lower computational requirements. While models such as RS-BYOL (Jain et al., 2022a) have demonstrated that such an approach is feasible with remotely sensed imagery, distillation has yet to be explored as a SSL approach in remote sensing to the degree that masked auto-encoding or contrastive learning with negative samples have been. The recent and well-publicized success of DeepSeek (Wu et al., 2024) in the NLP field highlights the viability of distillation as an SSL approach and indicates that it should be further examined within the remote sensing domain.

### v2

@@ 8.4. Multiple Decoders
The loss for all of these decoders is back propagated into the encoder, resulting in representations of optical imagery that capture data from other modes. While this approach learns refined optical representations, contrastive learning approaches of a similar manner would also learn representations for the other modes of data. Depending on the downstream modalities available, this may be a significant limitation for MMEarth’s effectiveness.
Feature Guided MAE (FG-MAE) (Wang et al., 2023b) also explores the utilization of multiple decoders for MAE, but does so by tasking the model to reconstruct specific features of interest in the remote sensing world. MAE’s focus on pixel-level details can limit the model’s capability in understanding satellite imagery. FG-MAE explores a variety of reconstruction objectives, including Histograms of Oriented Gradients (HOG) (Dalal and Triggs, 2005), Normalized Difference Indices (NDI) such as Normalized Difference Vegetation Index (NDVI) (Huang et al., 2021), CannyEdge (Canny, 1986), and Scale-Invariant Feature Transform (SIFT) (Lowe, 2004). FG-MAE maintains two separate models: one for optical imagery and one for SAR imagery. In EuroSat (Helber et al., 2019) and BigEarthNet (Sumbul et al., 2019) classification tasks, FG-MAE’s optical model demonstrates comparable or better results to standard MAE when only reconstructing a single one of these RS features. However, performing reconstruction on both HOG and NDI results in a substantial performance boost compared to standard MAE. FG-MAE’s SAR model finds that simply reconstructing HOG yields the best results, as it tends to work well with the noisier Sentinel-1 imagery.
@@ 8.5. Orientation Awareness
Masked Angle-Aware Auto-Encoding (M3AE) (Li et al., 2024b) emerges to address this particular problem. M3AE trains on the MillionAID dataset (Long et al., 2021), an RGB dataset derived from Google Earth imagery. M3AE introduces a scaling, center crop augmentation to every image, resulting in an augmented image where an object in the center will be rotated into a different orientation from the original. This augmented image is masked then fed into an encoder-decoder framework where the decoder attempts to reconstruct the original image. In order to ensure that the model pays attention to the re-oriented object, the encoder is fed positional embeddings that include a third dimension to track the angle of a given patch. M3AE also introduces an optimal transport reconstruction loss to automatically assign similar image patches for each rotated crop patch. This ensures that the rotation is taken into effect from a loss standpoint.
Rotated Varied-Size Attention (RVSA) (Wang et al., 2023c) also develops a novel approach to developing orientation-agnostic representations of RS imagery, but does so through ViT self-attention compared to image augmentation. Trained on the MillionAID dataset (Long et al., 2021), RVsA shifts away from the standard multi-head self-attention (MHSA) present in traditional CV ViTs, as full attention scales quadratically to image size. Given the potential high-resolution nature of RS imagery, RVSA leverages a varied-size window-based MHSA (VSA). In VSA, images are partitioned into non-overlapping windows, then have MHSA applied (Zhang et al., 2022). In order to learn orientation, RVSA randomly rotates and re-sizes these windows before performing self-attention within a ViT. With these self-attention tweaks in effect, standard MAE is employed for reconstruction masked MillionAID images. The result is a framework capable of swiftly handling high resolution RS imagery that learns orientation-agnostic representations of RS imagery.
@@ 8.6. Wavelength Awareness
Dynamic One-For-All (DOFA) (Xiong et al., 2024) attempts to address this problem by creating a framework that can take in data from any number of different sensors. DOFA follows a hypernetwork approach (Chauhan et al., 2024), where a hypernetwork takes in the wavelength details of a given band and then generates weights that are used to initialize a neural network. Image bands are masked, then fed into that neural network to generate encodings. The encodings are then fed through an encoder-decoder structure, with a mirrored hypernetwork at the end responsible for reconstructing the original image. This framework is designed to learn bandwidth specific features via the hypernetwork and shared features via the shared encoder-decoder. DOFA is trained on SAR imagery from Sentinel-1, multispectral imagery from Sentinel-2, hyperspectral imagery from enMAP, and high resolution RGB imagery from NAIP. The resulting framework achieves SOTA performance on downstream tasks such as BigEarthNet (Sumbul et al., 2019), So2Sat (Zhu et al., 2019), and EuroSat (Helber et al., 2019).
@@
Foundation Models for Forest Monitoring (FoMo) (Bountos et al., 2025) also trains to be aware of spectral wavelength, but does so without the use of a hypernetwork. FoMo trains on four satellite-based datasets and four aerial-based dataset, with spatial resolutions ranging from less than 5cm to 60m. The datasets include MS imagery from Landsat 8-9, Sentinel-2, and Gaofen-2. They also include SAR imagery from Sentinel-1, DEM data, and both RGB and MS imagery from high-res UAV sensors. The resulting training dataset is rich in spatial and spectral diversity. FoMo’s flexible training strategy groups together all of the images of the same spatiotemporal scene, then randomly selects bands to use in that particular sample. Each band is independently tokenized to avoid spatial resolution issues. FoMo then incorporates standard positional embeddings and introduces spectral embeddings to denote what band each channel belongs to. This setup allows for learned representations across bands and within bands. Tokens are masked, passed through a ViT encoder, and then go through a standard decoder for reconstruction. Due to the random band selection process for generating samples, batches can potentially be very different. Gradient accumulation across batches is used to ’smooth’ the learning rate.
FoMo’s architecture does not boast the ability to impute information for untrained on wavelengths the way DOFA might. However, the diversity of the training data and elegant approach yield a model that is effective at leveraging data from a wide swathe of sensors. The authors do note that independent channel tokenization can also lead to potential scalability issues. The amount of tokens per sample increases linearly with the number of spectral bands, resulting in the complexity of the self-attention mechanism increasing quadratically.
9. Combined Approaches
Several RS foundation models cannot in good conscience be described as falling into any of the above approaches. Instead, they leverage multiple objectives that span the approaches discussed above.
Contrastive Radar-Optical Masked Auto-encoders (CROMA) (Fuller et al., 2023) is one such model. CROMA trains on Sentinel-1 and Sentinel-2 imagery and trains an encoder for each sensor. However, CROMA takes inspiration from the Fast Language Image-Pretraining (FLIP) framework (Jia et al., 2021), which independently masks each mode of data before feeding them into encoders and performing contrastive learning in the standard CLIP (Radford et al., 2021) setup. Masking images before encoding them results in a significantly less memory intensive process than encoding the raw image when vision transformers are used. Therefore, the FLIP framework enables contrastive training with a much larger batch size with the same or less memory footprint. This results in faster convergence. CROMA implements this approach with remotely sensed images, contrastively learning between masked Sentinel-1 and Sentinel-2 images. CROMA also introduces a third encoder that takes in concatenated Sentinel-1 and Sentinel-2 encodings to generate multi-sensor encodings. These fused encodings are fed into a multi-sensor decoder responsible for constructing the initial Sentinel-1 and Sentinel-2 images. In this manner, CROMA is able to introduce contrastive and reconstruction loss terms into their framework. This results in a model that effectively learns uni-sensor representations for Sentinel-1 and Sentinel-2 and multi-sensor representations of both Sentinel-1 and Sentinel-2. CROMA’s framework is benchmarked against BigEarthNet (Sumbul et al., 2019) and FMoW (Christie et al., 2018) classification tasks, demonstrating SOTA performance with images from a single sensor and multi-sensed images. BigEarthNet contains roughly half a million Sentinel-2 images where each patch can have multiple land-cover classes. CROMA also utilizes modified Attention with Linear Biases (ALiBi) positional embeddings (Press et al., 2022) for their ViT encoders. These embeddings provide an advantage over standard sinusoidal positional embeddings by biasing the embeddings based on Euclidean distance between key-query pairs, thus imbuing the embeddings with a degree of spatial awareness.
OmniSat (Astruc et al., 2024) also leverages an encoder per modality, but does so for three separate modes of data:
timeseries of multispectral imagery
timeseries of SAR imagery
high resolution UAV imagery that features RGB bands, a NIR band, and DEM data
After patchifying and encoding the data through their respective encoders, OmniSat then performs contrastive loss between the patches of each modality, where each patch has M-1 positive pairs (M representing the number of modalities). However, OmniSat makes a key change to how it determines negative pairs, and denotes that any patches which neighbor the positive pair should not be considered as a negative pair. This may seem odd from the perspective of natural imagery, but within RS imagery, there tends to not be vastly different semantic information in neighboring patches. This follows Tobler’s first law of geography: ”Everything is related to everything else, but near things are more related than distant things” (Miller, 2004). For the purposes of contrastive learning, it means that trying to use those neighboring patches as negative pairs will likely result in false negatives. This contrastive strategy addresses that false negative issue. OmniSat then performs independent masking, combines the latent representations into a multi-modal representation, and then reconstructs the original image modes from the shared multi-modal representation. They choose MS timeseries reconstruction targets by considering the temporal attention maps. Lower self-attention scores indicate cloudy images, which make for poor reconstruction targets.
Cross-Scale MAE (Tang et al., 2024) also leverages both contrastive and MAE approaches. Where OmniSat and CROMA use this combined approach to better train on multi-sensor data, Cross-Scale MAE does so in order to better understand spatial resolution. Trained on RGB data from the FMoW (Christie et al., 2018) dataset, Cross-Scale MAE performs a random crop and resize augmentation on a sample, then compares that to the original image. This comparison contains three components:
contrastive loss between embeddings of the original image and augmented image
mean squared error between the output of the attention block of the decoder for the original and augmented images
reconstruction loss of the original image (from a patchified version of the original) and augmented image (from a patchified version of the augmented image)
Contrastive loss between different spatial resolutions of the same image ensures that the model treats images with the same subject matter the same, regardless of size. Mean Square Error (MSE) loss between the attention block outputs is designed to reinforce that relationship, and then standard reconstruction loss encourages learning intra-image semantics. OmniSat performs very well on the RESISC45 (Cheng et al., 2017) task, which features images with varying spatial resolution, compared to other spatially aware models, such as ScaleMAE (Reed et al., 2023). OmniSat explores adding the GSD embedding present in ScaleMAE and finds that it does not improve performance. The author’s theorize that it provides redundant information, and further point to the fact that incorporating GSD on SatMAE (Cong et al., 2023), which has no spatial awareness, improves performance.
Geospatial Foundation Model (GFM) (Mendieta et al., 2023) also leverages multiple approaches to train on a single sensor. Where Cross-Scale MAE focuses on understanding spatial resolution, GFM focuses on treating RS imagery as an extension of natural imagery, performing continual pre-training with models trained on the ubiquitious ImageNet (Russakovsky et al., 2015) dataset. GFM draws inspiration from the fact that that MIM can trivialize the task when training on data with low entropy. In other words, if there is not much difference from patch to patch, then the network simply does not work as hard to reconstruct a masked patch. GFM calculates the average entropy of 3k randomly selected images from ImageNet and Sentinel-2 and find that Sentinel-2 images have an entropy of 3.9, compared to an entropy of 5.1 for ImageNet (Mendieta et al., 2023). GFM develops a two-pronged learning approach to address this issue. GFM maintains a student encoder and teacher encoder, but the teacher encoder is an entirely frozen model initialized with weights from training on ImageNet. For each image, the student is fed a masked version of the image, and the teacher is fed an unmasked version. Then, there are two loss terms generated:
the cosine similarity between the teacher and the student embeddings
reconstruction loss from trying to reconstruct the original image through the student embeddings
This framework is designed to leverage pre-exisiting work in CV via distillation to minimize the overhead needed to train RS foundation models. Distilling against a pre-trained ImageNet model lets GFM quickly learn how to extract features from images, taking advantage of the higher entropy present in the ImageNet images. The additional MIM task ensures that GFM learns RS domain-specific information.
@@ 10.1. Learning Representations in the Era of Climate Change
Given the rapidly shifting nature of the Earth’s climate, learned representations should be robust enough to take these changes into account without substantial retraining. Otherwise, the re-usability of these representations becomes less desirable and/or reliable over time, diminishing one of the main benefits of foundation models. When considering the variety of modalities present in RS, there is the potential to create representations that can withstand the challenges of a changing climate, but a degree of care is required. Relying too heavily on geographic locations and temporal data could very well result in representations that quickly become outdated as the world’s climate shifts. For example, a learned representation of a specific location on a coastline could cease to be relevant due to rising sea levels. Similarly, climate change has resulted in marked seasonal shifts within the past decade. Therefore, learning representations of what a given location looks like at a given time of year may wane in relevance as well.
By comparison, representations based on modalities that directly observe the Earth, such as satellite imagery and text, should be more tolerant to climate change-induced drift. Such representations are tied to understanding underlying observations directly within the data and are less reliant on climate context. Given that many remotely sensed datasets date back decades, there is merit in creating foundation models that perform time-series forecasting over long time periods in order to effectively learn representations which take long-term trends into account. This historical data also opens the door for assessing the capability of RS foundation models to quantify historically understood climate and biome changes.
The combination of historical data and the steady incoming stream of satellite data from currently orbiting satellite offers intriguing possibilities for continual pre-training, an approach popularized in NLP (Yıldız et al., 2025). Continual pre-training focuses on efficiently adapting existing models to emerging knowledge in a field while maintaining previously learned knowledge. Within the remote sensing domain, GFM (Mendieta et al., 2023) has begun to explore this approach. However, GFM does so by continually training a CV model against RS data, instead of continually training an existing RS model against a stream of new RS data. Developing a foundation model on historical data, and then performing continual pre-training against up-to-date RS data could prove useful for the development of climate change-tolerant RS foundation models. This approach may prove particularly pertinent when working with spatiotemporally dependent data modes, where the same scene may be re-contextualized due to climate change.
@@ 10.2. Under-Tapped Modes for Geospatial Representation Learning
Much of the current research focuses on drawing parallels to the work done on CV foundation models. Given the success of foundation models in the CV domain and the sheer amount of imagery that is available, be it from satellites or other aerial sources, this is a necessary first step. However, models such as MMCPP (Chen et al., 2023) use the CLIP contrastive learning objective and rely on other modalities, such as trajectory maps, points of interest, and aerial images. Their success should serve as a somewhat cautionary tale against becoming too invested in foundation models that purely focus on satellite and/or aerial imagery. Environmental stations measure features such as groundwater, weather, or snow melt data. They provide important climate and atmospheric information at a very high temporal resolution, often recording data on a daily or even hourly basis. More human-centric data, such as social media posts, social connectedness indices (Bailey et al., 2018), and trajectory data, also provide invaluable information for understanding more societally-focused downstream tasks. Both forms of information are not observable from space, but could provide useful modes due to their different perspective and high temporal resolution. Data on rapidly changing phenomena also tends to not be observable from space. Changes can occur too quickly for a satellite to capture due to orbit and revisit time. GAIR’s (Liu et al., 2025) usage of INR representations for satellite imagery also provides a useful starting point for mapping granular, ground-level data to satellite imagery.
@@ 10.3. Performance versus Training Expenses
Given the computational costs of training transformer models, the training expense of foundation models has rapidly become an obstacle across a variety of domains, with some models now including a section on carbon footprint (Cong et al., 2023). As a substantial portion of downstream tasks for remote sensing pertain to combating the negative effects of climate change, there is a compelling argument for remote sensing foundation models to focus on curtailing training expenses. Massive models which create a large carbon footprint ultimately may be counterproductive from a climate change standpoint. However, the perpetual flow of remotely sensed data which dates back decades also encourages exploring the development of an incremental learning pipeline to efficiently fine-tune foundation models with up-to-date information.
@@
Aside from a carbon footprint perspective, models with huge numbers of parameters have their downstream applications somewhat limited due to computational requirements. Not all users have the hardware necessary to fine-tune foundation models with a high parameter count. Some models, such as MOSAIKS (Rolf et al., 2021), focus on lightweight frameworks in order to promote access to remote sensing foundation models. MOSAIKS has competitive results on a number of downstream tasks, indicating that increasing parameters within RS foundation models may follow a law of diminishing returns. On the other hand, the encoder-decoder structure that is predominant in foundation models allows for repeated, diverse usage of representations on downstream tasks. Arguably, leveraging significant resources to generate high quality embeddings as a one-time cost is the most effective path moving forward, as this approach results in SOTA performance on downstream tasks with relatively few resources necessary for future fine-tuning.
Other domains have specifically studied how the factors of data size, number of parameters, and compute affect overall model performance (Kaplan et al., 2020). NLP has established a principle that performance increases with an increase in any of those factors, providing that neither of the other two serve as a bottleneck (Kaplan et al., 2020). However, such studies in CV for MIM have demonstrated that large-scale data is a requirement for scaling up compute and model parameters (Xie et al., 2023). Moreover, MIM cannot benefit from increasing the data size if the model is not overfitting (Xie et al., 2023). To the best of our knowledge, such studies do not exist for the remote sensing field. Authors in CV cite the semantic richness in text tokens compared to the redundancy of visual tokens as a significant motivator for their study (Xie et al., 2023). It stands to reason that the significant differences in natural imagery and remotely sensed imagery would also motivate such a study in the remote sensing field.
@@
It is possible that SSL distillation networks may provide a happy medium between the desire for high quality embeddings and lower computational requirements. While models such as RS-BYOL (Jain et al., 2022a), DINO-MM (Assran et al., 2023), and SkySense (Guo et al., 2024) have demonstrated that such an approach is feasible with remotely sensed imagery, distillation has yet to be explored as a SSL approach in remote sensing to the degree that masked image modelling or contrastive learning with negative samples have been. The recent and well-publicized success of DeepSeek (Wu et al., 2024) in the NLP field also highlights the viability of distillation for knowledge compression and indicates that it should be further examined within the remote sensing domain. GFM (Mendieta et al., 2023) begins to inspect the efficacy of distilling knowledge from the CV domain into remote sensing, but there remains substantial room for exploration.
A Mixture of Experts (MoE) approach may also provide an intriguing compromise between embedding quality and computational requirements. At a high level, MoE partitions an input space into non-overlapping regions, then designates an ’expert’ model per region (Gamal et al., 2025). This particular approach allows for flexible incorporation of models for a variety of modalities and tasks, yet remains computationally efficient since only a subset of experts are activated at training and inference. The Switch Transformer (Fedus et al., 2022), an NLP model for multilingual learning, followed the MoE paradigm to develop a foundation model with over a trillion trainable parameters. In addition to the enormous parameter count, the Switch Transformer also trains significantly faster than comparable models on the same hardware (Fedus et al., 2022). The MoE framework’s capabilities in efficiently leveraging mode and task specific models make it an exciting avenue for exploration in RS foundation models. To the best of our knowledge, a MoE-based remote sensing foundation model does not yet exist.